OpenStudy (anonymous):

Is the definite integral of this 0? (π/4) ∫ sec^(2)(x)x (0)

OpenStudy (anonymous):

I got 0 but im not sure if I'm right..

OpenStudy (anonymous):

I got 1. Verify that your integrating sec(x)^2

OpenStudy (anonymous):

=[tanx]from 0 to pi/4=1

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4}\sec(x)^2 dx = 1\]